A person with his ear to the ground, sees a huge stone strikethe concrete pavement. A moment later two sounds are heard from theimpact: one travels n the air and the other in the concrete, andthey are 2.2 s apart. How far away did the impact occur?
*The speed of sound for air is 343 m/s, for concrete˜3000.
Please show how you set up the problem and how you got youranswer.
Answer #1
This is is really quite simple.
velocity*time=distance
vt=d
the distance is the same for both sounds.
3000t=d=343(t+2.2)
3000t=343(t+2.2)
solving for time, we get t=.284 seconds.
now to find the distance, we just plug this into the firstequation.
(3000m/s)*(0.284s) = 852 meters.